OpenStudy (anonymous):

Show that \[\int\limits\limits {1-(\tan \theta)^2 \over [1 +(\tan \theta)^2]^2}theta = \int\limits \cos2 \theta d \theta\]

OpenStudy (lgbasallote):

no offense..but i really find the inconsistency of your math topics funny =)))) anyway...for this question uhmmm...

OpenStudy (anonymous):

dx is on the left dtheta on the right?

OpenStudy (lgbasallote):

hmm good point -_-

OpenStudy (anonymous):

Yes, it is

OpenStudy (lgbasallote):

maybe \(1 + \tan^2 \theta = \sec^2 \theta\) will be useful here??

OpenStudy (anonymous):

so on the left side, you're integrating wrt theta, and there's no theta... are those theta's supposed t be x's?

OpenStudy (shubhamsrg):

miss emma,,that cant be ..please recheck the ques..

OpenStudy (shubhamsrg):

i mean both dx and dtheta cant be there

OpenStudy (lgbasallote):

i think she already corrected herself that it's \(d\theta\) :p

OpenStudy (anonymous):

It's meant to come out as theta but it starts of as dx

OpenStudy (lgbasallote):

wait..what?

OpenStudy (anonymous):

It's not yet integrating.. first it's identities...

OpenStudy (anonymous):

That's what the question says

OpenStudy (lgbasallote):

so...we need to change it to x...without knowing the relationship of theta to x?

OpenStudy (shubhamsrg):

well in that case ,,if we assume ques to be correct,,we should suppose theta is constant with dx ..and then simplify,,which is much easier.. and the solution obtained should be integral of cos2theta dtheta

OpenStudy (anonymous):

Yup, but how to get there? I don't know... I'm not good with identities and trig.

OpenStudy (shubhamsrg):

well then,, multiply and divide LHS by \[\cos ^{4}\theta \] convert to sin and cos terms..

OpenStudy (shubhamsrg):

the denominator then becomes 1..

OpenStudy (shubhamsrg):

numerator is cos 2(theta) * cos^2 (theta)

OpenStudy (shubhamsrg):

which is definately not the integral of cos2(theta) so please re check the ques,,else i'd assume ques to be incorrect..

OpenStudy (anonymous):

OH, yes sorry.. it was first dx, and then I changed it myself to theta, so it should be theta.

OpenStudy (shubhamsrg):

i still think its incorrect

OpenStudy (anonymous):

It's theta to theta...

OpenStudy (anonymous):

That's the question

OpenStudy (shubhamsrg):

i re checked my soln and the ques is wrong..

OpenStudy (anonymous):

That\s the question here...

OpenStudy (anonymous):

It has something to do with dx = sec2^θ dθ

OpenStudy (shubhamsrg):

theres a cos^2 (theta) extra coming in the numerator on LHS along with cos 2(theta)

OpenStudy (shubhamsrg):

dx=sec^2 theta dtheta ?? whats this?? you derived this ?

OpenStudy (anonymous):

In the answer, it says State or imply dx = sec^2θ dθ

OpenStudy (shubhamsrg):

yes,,theere we go..in presence os sec^2 theta,,cos^2 theta gets cancelled and we get our answer,,which makes it correct miss watson..

OpenStudy (anonymous):

Can you show via latex? It's easier to read and understand for me :D

OpenStudy (shubhamsrg):

\[1−(\tanθ)^2[1+(\tanθ)^2]^2 \] multiply and divide this by (cos(theta))^4 reduce expression into purely sin and cos terms.. the denominator becomes 1 as it is (sin^2(theta) + cos^2(theta))^2 now num is cos^2(theta) * cos (2theta) since you say dx = sec^2(theta),, we are left with only cos(2theta) hope it was clear now?

OpenStudy (lalaly):

is the fraction multiplied by theta? Can u rewrite ur problem ?

OpenStudy (lalaly):

When u decide to fix this tag me

OpenStudy (lalaly):

The right identity is\[\large{\frac{1-\tan^2x}{1+\tan^2x}=\cos(2x)}\]

OpenStudy (experimentx):

that makes sense!!